20% of the Chocos packets contain a free toy. Whats the probability youll need to open fewer than 4 packets before finding your first toy?


1
Expert's answer
2022-05-31T15:34:52-0400

The formula for geometric distribution is


"P(X=x)=(1-p)^xp"

Given "p=0.2"

"P(X<4)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"=(1-0.2)^0(0.2)+(1-0.2)^1(0.2)"

"+(1-0.2)^2(0.2)+(1-0.2)^3(0.2)"

"=0.2+0.16+0.128+0.1024=0.5904"